SOME HISTORY OF THE CONJUGATE GRADIENT AND LANCZOS ALGORITHMS: 1948–1976

Gene H. Golub and D. P. O’Leary

Received September 1, 2000

Abstract. This paper gives some of the history of the conjugate gradient and Lanczos algorithm and an annoteated bibliography for the period 1948–1976.

1 Algebra This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics.

2000 Mathematics Subject Classification. Primary 65F10, 65F15; Secondary 65H10, 65F03.

Key words and phrases. conjugate gradient algorithm, Lanczos algorithm, variable metric algorithm.

2 Geometry This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics.

3 Analysis This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics. This is a paper on mathematics.